OpenStudy (cutiekawdfi):

elp please ((: fan and medal <3 1. multiply and simplify if possible 4 √11 times 4 √10 2. what is the simplest form of the expression 3 √24a^10b^6 3. what is the simplest form of the quotient 3 √4x^2 times 3 √8x^7

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[4\sqrt{11}\times 4\sqrt{10}\] or \[\sqrt[4]{11}\times \sqrt[4]{10}\]?

OpenStudy (misty1212):

if it the first one, since \(4\times 4=16\) and \(10\times 11=110\) it would be \[16\sqrt{110}\]

OpenStudy (misty1212):

if it is the second one, then just \[\sqrt[4]{110}\]

candycove (candycove):

You want elp?

OpenStudy (anonymous):

Lmao

OpenStudy (cutiekawdfi):

-_-

OpenStudy (cutiekawdfi):

rude retricemother flutterer